1. Uranium-238 decays by a number of steps of alpha and beta decays to Lead-206. How many alpha paricles and how many beta particles are emitted?

2. Calculate the wavelength in nanometers of the blue light emitted by a mercury lamp with a frequency of 853.6 MHz.

3. Calculate the energy per mole of photons with a wavelength of 578 nm.
Answer #1

2)

nm of light

f = 853.6 MHz = 853.6*10^6 Hz = 8.536*10^8 Hz

WL = c/f = (3*10^8) / (8.536*10^8 ) = 0.351452 m

WL = 0.351452 m = 0.351452 *10^-9 nm

NOTE that the actual value of blue light is around 700 nm

3)

WL = 578 nm = 578*10^-9 m

E = hc/WL

E = (6.606*10^-34)*(3*10^8)/(578*10^-9 ) =3.428*10^-19 J/photon

E= 3.428*10^-19 J/photon *6.022X10^23 photon / mol phton = 206434.16 J/mol photon